# From Heisenberg and Schrödinger to the $\mathsf{P}$ vs. $\mathsf{NP}$ Problem

### v1

@@ 2.1.1 Bohr-Sommerfeld
with corresponding momentum:
$p(t)=m\dot{x}(t)=im\sum_{\alpha}\omega_{\alpha}a_{\alpha}\,e^{i\omega_{\alpha}t}.$ (5)
@@
From the expressions (4), (5), and (7), the action variable can be obtained as:
$J=\oint p\,dq=2\pi m\sum_{\alpha}\alpha\,\omega_{\alpha}\,|a_{\alpha}|^{2},$ (9)
which follows from integrating $p(t)\dot{x}(t)$ over one period of the motion.
In the old quantum theory, this action variable was constrained by the quantization rule (7), thereby selecting discrete orbits from the continuous set of classical trajectories.
@@ 2.1.2 Heisenberg
$\omega(n,n-\alpha)=\frac{E_{n}-E_{n-\alpha}}{h},$ (11)
@@ 2.1.5 Dirac
$\dot{x}=[x,H]\equiv xH-Hx,$ (36)
@@ 2.1.6 Heisenberg
In 1927, Heisenberg published his seminal paper “Über den anschaulichen Inhalt der quantentheoretischen Kinematik und Mechanik” on the uncertainty principle [20]. In this work, he cited Max Born for their joint development of matrix mechanics with Pascual Jordan and credited Jordan specifically for the Dirac–Jordan transformation theory. He acknowledged Wolfgang Pauli for his influence, discussions, and early formulation of related ideas, thanking him explicitly in a footnote. Paul Dirac is mentioned repeatedly, particularly in connection with transformation theory and the commutation relations, as Heisenberg directly linked his own results to Dirac’s formalism. Heisenberg treats Paul Ehrenfest’s work as foundational and insightful, building upon it. Louis de Broglie appears in a positive light, associated with wave–particle duality and the role of de Broglie wavelengths in defining the limits of spatial and temporal measurements. Albert Einstein is cited respectfully for the Compton effect, the photoelectric effect, and his insights into the wave–particle relation. Niels Bohr, mentioned in the footnotes as a source of inspiration, is credited for formative conversations on the limits of classical visualization; Heisenberg explicitly notes that the paper evolved from ideas developed in collaboration with Bohr [20].
@@ 2.2.1 De Broglie and Einstein
$\nu=\frac{m_{0}c^{2}}{h\sqrt{1-v^{2}/c^{2}}}.$ (46)
@@ 2.2.2 Schrödinger
$\frac{p}{\hbar}=\nabla S=k,\qquad\frac{E}{\hbar}\,=-\frac{\partial S}{\partial t}=\omega.$ (57)
@@ 2. Gradient:
$\nabla\psi=\frac{i}{\hbar}(\nabla S)\,k\,\psi.$ (61)
@@ 2.2.3 Ehrenfest
$\frac{d}{dt}\langle A\rangle=\frac{i}{\hbar},\langle[H,A]\rangle\left\langle\frac{\partial A}{\partial t}\right\rangle.$ (93)
@@ 2.2.4 Dirac
$\big(c\,\langle B|\big)|A\rangle$ $=c\,\langle B|A\rangle.$ (100)
@@ 2.2.6 Born
Born recognized that these amplitudes cannot be interpreted causally. Quantum mechanics does not predict the specific outcome of an individual collision but only the likelihood of possible outcomes. He therefore proposed that the probability for the transition $n\!\rightarrow\!m$ into a given solid angle is proportional to the modulus squared of the amplitude [5, 6]:
@@ 2.3. Hilbert Space
Although the demonstrations by Schrödinger, Pauli, and others showed striking agreement between matrix and wave mechanics in specific solvable systems—such as the harmonic oscillator and the hydrogen atom—this did not amount to a general proof of equivalence. As Fred Muller has demonstrated in his detailed historical–philosophical analysis, the 1926–1927 arguments were not mathematically conclusive, nor could they have been at the time. The two formulations remained only partially aligned, differing in mathematical structure, empirical scope, and ontological commitments. A fully rigorous equivalence emerged only later, with John von Neumann’s 1932 Hilbert-space formulation, which unified them within a common state–observable framework. Until then, the “equivalence” remained, in Muller’s words, “a historically contingent myth” [24].
The search for a truly unified and mathematically rigorous framework did not end with Schrödinger and Dirac. It reached its culmination in the work of von Neumann, whose formulation of quantum mechanics between 1927 and 1932 transformed the heuristic operator calculus of physicists into the abstract and complete structure of Hilbert space.
Dirac’s The Principles of Quantum Mechanics (1930, first edition) contains no mention of von Neumann, even though von Neumann had already published his 1927 paper and several follow-up studies leading to his 1932 book. This omission was not accidental but reflected a difference in method and purpose. Dirac and von Neumann approached the foundations of quantum mechanics from complementary directions: Dirac developed a symbolic algebra of states and observables (bras, kets, $\delta$-functions, completeness) based on physical intuition and continuous transformations, whereas von Neumann built a rigorous mathematical structure grounded in measure theory and functional analysis.
Von Neumann’s work unified the contributions of Born, Heisenberg, Jordan, Schrödinger, and Dirac into a single coherent framework. He demonstrated that the matrix and wave formulations are not different theories but different representations of the same underlying Hilbert-space structure. Schrödinger’s differential operators and Heisenberg’s infinite matrices are merely coordinate descriptions of self-adjoint operators acting on this space. Von Neumann’s 1927–1932 papers and book constitute the first mathematically rigorous unification of all existing forms of quantum mechanics within one abstract framework.
Von Neumann begins his 1927 paper, “Mathematische Begründung der Quantenmechanik,” by explicitly observing that both approaches — matrix mechanics and Schrödinger’s differential equations acting on continuous wave functions — describe the same eigenvalue problem in different forms. He writes that in the matrix method one deals with ”unendliche Matrizen” (infinite matrices), while in the differential method one deals with functions $\psi(q)$ defined over a domain $\Omega$. He then constructs a common space containing both [33]:
@@
showing that both are manifestations of one mathematical object, the Hilbert space [33].
This move was conceptually revolutionary because it transformed the physicists’ patchwork of infinite-dimensional symbols into a precise linear space with an inner product, enabling the rigorous definition of orthogonality, completeness, and convergence.
Then, von Neumann introduced self-adjointness (Selbstadjungiertheit) as the precise mathematical criterion for an operator $H$ to represent an observable. By formulating the correct boundary conditions, he ensured that self-adjoint operators yield real eigenvalues and generate unitary time evolution [33]. This step transformed the physicists’ informal use of “Hermitian” matrices and differential operators with real eigenvalues into a rigorous concept—one that became the cornerstone of functional analysis in quantum mechanics.
Afterward, von Neumann introduces the term ”unitär” (unitary) for transformations preserving the inner product [33]:
@@
This provides the first rigorous mathematical definition of the unitary transformation, serving as the bridge between representations (matrix $\leftrightarrow$ wave $\leftrightarrow$ momentum). This was a direct formalization of Dirac’s and Jordan’s [8] transformation theory, but now without any of their $\delta$-function heuristics or ambiguous probability amplitudes.
Von Neumann devotes a section to criticizing Dirac’s use of improper functions such as $\delta(x-x^{\prime})$ and non-normalizable eigenfunctions. He shows that these are mathematically inconsistent within the existing framework and argues that one must instead use spectral measures on Hilbert space—an idea he develops fully in forthcoming publications and his 1932 book Mathematische Grundlagen der Quantenmechanik [33, 34]. So the novelty here is not yet the final spectral theorem but the recognition that the continuous spectrum can be handled rigorously without $\delta$-functions, by redefining what eigenfunctions mean.
He rephrases both matrix and differential eigenvalue equations as a single operator equation $Hf=af$, where $f$ belongs to a Hilbert space $\Phi$, which may consist of sequences $x_{1},x_{2},\ldots$ or functions $\psi(q)$ satisfying [33]:
$\sum_{i}|x_{i}|^{2}<\infty\quad\text{or}\quad\int|\psi|^{2}\,dq<\infty.$ (120)
This abstraction marks the birth of the modern spectral approach, the idea that all quantum observables correspond to linear operators on an abstract state space.
Physicists such as Born, Jordan, and Dirac had reached the correct physical conclusions (probabilistic interpretation, amplitudes, unitarity), but their mathematics was loose. Von Neumann’s 1927 paper—and several that followed—were novel in kind rather than in content, recasting the entire quantum theory into a framework of rigorous mathematical analysis. He proves convergence, defines the domains of operators, and requires that the scalar products and norms be finite, thereby establishing the mathematical discipline that would dominate all later quantum theory.
Later, von Neumann treated states as quadratic forms and observables as operators. His 1927–1932 work introduced the density matrix, completing Born’s statistical rule in an operator language. Already in his first paper, he distinguishes the state vector as a point in Hilbert space; the observable as a self-adjoint operator acting on it; and the probability interpretation as emerging from inner products. Von Neumann transformed the heuristic operator calculus of early quantum mechanics into a rigorous theory of states and observables formulated within the framework of Hilbert space.
In his 1932 book, Mathematical Foundations of Quantum Mechanics, von Neumann gives a rigorous unification of wave and matrix mechanics by identifying both with the same abstract Hilbert space framework. He begins from the Riesz–Fischer isomorphism between the sequence space of matrix mechanics [34]:
$F_{Z}=\Bigl\{\,x=(x_{\nu})\;\Big|\;\sum_{\nu=1}^{\infty}|x_{\nu}|^{2}<\infty\Bigr\},$ (121)
@@
$F_{\Omega}=L^{2}(\Omega)=\Bigl\{\,\phi:\Omega\to\mathbb{C}\;\Big|\;\int_{\Omega}|\phi(q)|^{2}\,dq<\infty\Bigr\},$ (122)
which are linearly isometric via an inner-product preserving correspondence:
@@
Under this isometry, the basic observables match: position and momentum in the wave picture, $q_{j}$ (multiplication by $q_{j}$) and ${\dfrac{\hbar}{i}}{\dfrac{\partial}{\partial q_{j}}}$, correspond to matrices $Q_{j},P_{j}$ in the sequence picture. Both realizations satisfy the canonical commutation relations [34]:
@@
so any Hamiltonian $H$ formed from these by the same algebraic operations yields identical numerical predictions in either representation.
Von Neumann then abstracts from these concrete models to define Hilbert space as a complete, separable complex inner-product space. He treats states as (unit) vectors and develops the geometry of orthonormal sets, expansions:
@@
and completeness. He next systematizes operators on Hilbert space, i.e., domains, adjoints $A^{\ast}$, self-adjoint (Hermitian) and unitary operators, and the distinction between bounded and unbounded operators.
The stationary Schrödinger problem becomes the abstract eigenvalue problem $H\phi=\lambda\phi$ for a self-adjoint $H$. Von Neumann formulated probabilities as quadratic forms on Hilbert space [34]:
@@
where $(\phi,\psi)$ denotes the Hilbert-space inner product. $E_{A}(\Delta)$ is the projection operator associated with the spectral measure of $A$. In a particular representation, for instance, when $A=Q$ acts by multiplication on $L^{2}(\mathbb{R}^{3})$, this formula reduces to Born’s familiar rule:
@@
Thus, the abstract expression $\psi,E_{A}(\Delta)\psi$ provides the rigorous, representation-independent foundation of Born’s probability law.
@@
In this way, von Neumann completed the transition from quantum mechanics as a calculational scheme to a fully axiomatized physical theory, in which states, observables, and probabilities acquire precise mathematical meaning within a single abstract space. The upshot is that matrix mechanics and wave mechanics are two unitarily equivalent representations of a single theory whose states are Hilbert-space vectors (or density operators) and whose observables are self-adjoint operators satisfying the canonical commutation relations.

### v2

@@ 2.1.1 Bohr-Sommerfeld
with the velocity of the electron in its classical periodic orbit:
$\dot{x}(t)=\sum_{\alpha=-\infty}^{\infty}i\,\omega_{\alpha}(n)\,a_{\alpha}(n)\,e^{i\omega_{\alpha}(n)\,t},$ (5)
and with corresponding momentum:
$p(t)=m\dot{x}(t)=im\sum_{\alpha=-\infty}^{\infty}\omega_{\alpha}(n)\,a_{\alpha}(n)\,e^{i\omega_{\alpha}(n)t}.$ (6)
@@
From the expressions (4), (5), (6), and (8), the action for this orbit is:
$J(n)=\oint p\,dq=\int_{0}^{T(n)}p(t)\,\dot{x}(t)\,dt=m\int_{0}^{T(n)}\dot{x}(t)^{2}\,dt.$ (10)
The action variable can be obtained as:
$J(n)=2\pi m\sum_{\alpha=-\infty}^{\infty}\alpha\,\omega_{\alpha}(n)\,|a_{\alpha}(n)|^{2},$ (11)
where the sum runs over all integer harmonics $\alpha$.
In the old quantum theory, this action variable was constrained by the quantization rule (8), thereby selecting discrete orbits from the continuous set of classical trajectories. Thus, only those periodic trajectories satisfying $J(n)=nh$ were allowed.
@@ 2.1.2 Heisenberg
$\omega(n,n-\alpha)=\frac{E_{n}-E_{n-\alpha}}{\hbar}.$ (13)
@@ 2.1.5 Dirac
$\dot{x}=\frac{i}{\hbar}[H,x]\equiv\frac{i}{\hbar}(Hx-xH),$ (38)
@@ 2.1.6 Heisenberg
In 1927, Heisenberg published his seminal paper “Über den anschaulichen Inhalt der quantentheoretischen Kinematik und Mechanik” on the uncertainty principle [20]. In this work, he cited Max Born for their joint development of matrix mechanics with Pascual Jordan and credited Jordan specifically for the Dirac–Jordan transformation theory. He acknowledged Wolfgang Pauli for his influence, discussions, and early formulation of related ideas, thanking him explicitly in a footnote. Paul Dirac is mentioned repeatedly, particularly in connection with transformation theory and the commutation relations, as Heisenberg directly linked his own results to Dirac’s formalism. Heisenberg treats Paul Ehrenfest’s work as foundational and insightful, building upon it. Louis de Broglie appears in a positive light, associated with wave–particle duality and the role of de Broglie wavelengths in defining the limits of spatial and temporal measurements. Albert Einstein is cited respectfully for his contributions to the photoelectric effect and his insights into the wave-particle duality. Niels Bohr, mentioned in the footnotes as a source of inspiration, is credited for formative conversations on the limits of classical visualization; Heisenberg explicitly notes that the paper evolved from ideas developed in collaboration with Bohr [20].
@@
Heisenberg introduced the $\gamma$-microscope as a thought experiment intended to illustrate the reciprocal limitations involved in measuring the position and momentum of an electron. The idea is that to localize an electron very precisely, one must illuminate it with radiation of extremely short wavelength, such as $\gamma$-rays. Short wavelength improves the spatial resolution $\Delta x$, but $\gamma$-photons carry large momentum; when a photon is scattered into the microscope, it imparts an unpredictable momentum kick to the electron. Thus, decreasing $\Delta x$ by using shorter wavelengths necessarily increases the disturbance in momentum $\Delta p$. Heisenberg used this to motivate the uncertainty relation. In his 13 July 1935 letter to Einstein, Schrödinger criticizes this device, arguing that a $\gamma$-microscope does not in fact measure any variable that possesses a true canonical conjugate partner in the operator sense of equation (44) [35] (see the discussion in section 2.3.2).
@@ 2.2.1 De Broglie and Einstein
$\nu=\frac{\gamma m_{0}c^{2}}{h},\qquad\gamma=\frac{1}{\sqrt{1-v^{2}/c^{2}}}.$ (48)
@@ 2.2.2 Schrödinger
$\nabla S=p,\qquad-\,\frac{\partial S}{\partial t}=E,\qquad k=\frac{p}{\hbar},\qquad\omega=\frac{E}{\hbar}.$ (59)
@@ 2. Gradient:
$\nabla\psi=\frac{i}{\hbar}(\nabla S)\psi\,=\,ik\psi.$ (63)
@@ 2.2.3 Ehrenfest
$\frac{d}{dt}\langle A\rangle=\frac{i}{\hbar}\,\langle[H,A]\rangle+\left\langle\frac{\partial A}{\partial t}\right\rangle.$ (95)
@@ 2.2.4 Dirac
$\big(c\,\langle B|\big)|A\rangle$ $=c^{*}\,\langle B|A\rangle.$ (102)
@@ 2.2.6 Born
We compare Born’s exact asymptotic scattering form:
$\Psi(\mathbf{r},q_{k})\sim e^{2\pi i\,\mathbf{k}\cdot\mathbf{r}}+\sum_{m}f_{nm}(\hat{\mathbf{r}})\,\frac{e^{ikr}}{r}\,\psi_{m}(q_{k}),$ (115)
with the schematic expression (114). The two expressions are equivalent once we identify the incoming wave in (114) with the incident plane wave of (115):
$\text{(incoming wave)}\equiv e^{2\pi i\,\mathbf{k}\cdot\mathbf{r}},$ (116)
[…]
$\Phi_{nm}(\hat{\mathbf{r}})\equiv f_{nm}(\hat{\mathbf{r}})\,\frac{e^{ikr}}{r},$ (117)
where $\dfrac{e^{ikr}}{r}$ is the standard outgoing spherical-wave factor in the asymptotic form of the scattering solution.
Born recognized that the scattering amplitudes cannot be interpreted causally. Quantum mechanics does not predict the specific outcome of an individual collision but only the likelihood of possible outcomes. He therefore proposed that the probability for the transition $n\!\rightarrow\!m$ into a given solid angle is proportional to the modulus squared of the amplitude [5, 6]:
@@ 2.3. Hilbert Space
2.3.1 Von Neumann
Although the demonstrations by Schrödinger, Pauli, and others showed striking agreement between matrix and wave mechanics in specific solvable systems—such as the harmonic oscillator and the hydrogen atom—this did not amount to a general proof of equivalence. As Fred Muller has demonstrated in his historical–philosophical analysis, the 1926–1927 arguments were not mathematically conclusive, nor could they have been. The two formulations remained only partially aligned, differing in mathematical structure, empirical scope, and ontological commitments. A fully rigorous equivalence emerged only later, with John von Neumann’s 1932 Hilbert-space formulation, which unified them within a single state–observable framework. Until then, the “equivalence” remained, in Muller’s words, “a historically contingent myth” [24].
The search for a truly unified and mathematically rigorous foundation did not end with Schrödinger and Dirac. It reached its culmination in the work of von Neumann, whose contributions between 1927 and 1932 transformed the heuristic operator calculus of early quantum mechanics into the abstract and complete structure of Hilbert space. Dirac’s Principles of Quantum Mechanics (1930) contains no reference to von Neumann, even though von Neumann had already published his 1927 paper and several follow-up studies. The omission reflects a difference in method and purpose: Dirac developed a symbolic calculus of states and observables based on physical intuition, whereas von Neumann built a rigorous mathematical theory grounded in measure theory and functional analysis.
Von Neumann unified the contributions of Born, Heisenberg, Jordan, Schrödinger, and Dirac into one coherent theoretical structure. He showed that matrix and wave mechanics are not different theories but different representations of the same abstract Hilbert-space framework. Schrödinger’s differential operators and Heisenberg’s infinite matrices become coordinate descriptions of self-adjoint operators acting on that space. Von Neumann’s 1927–1932 papers and his 1932 book constitute the first mathematically rigorous unification of all existing forms of quantum mechanics.
In his 1927 paper, “Mathematische Begründung der Quantenmechanik,” von Neumann begins by observing explicitly that matrix mechanics and Schrödinger’s differential method both address the same eigenvalue problem in different languages. The matrix method employs unendliche Matrizen (infinite matrices), while wave mechanics uses functions $\psi(q)$ on a domain $\Omega$. He then constructs a common space containing both [36]:
@@ 2.3.1 Von Neumann
showing that both are instances of one mathematical entity, Der Hilbertsche Raum, the Hilbert space. This move was conceptually novel, transforming a patchwork of infinite-dimensional symbols into a precise linear space with an inner product, in which orthogonality, completeness, and convergence are rigorously defined.
Von Neumann next introduced self-adjointness (Selbstadjungiertheit) as the mathematically correct criterion for an operator to represent an observable. By specifying proper domains and boundary conditions, he ensured that self-adjoint operators have real spectra and generate unitary time evolution [36]. This replaced the physicists’ informal use of Hermitian operators with a rigorous concept that became the cornerstone of functional analysis in quantum theory.
He then defined unitär (unitary) transformations as those preserving the inner product [36]:
@@
This is the first mathematically precise definition of a unitary operator, formalizing Dirac’s and Jordan’s transformation theory [8] without relying on $\delta$-function heuristics or ambiguous probability amplitudes.
Von Neumann also criticized Dirac’s use of improper eigenfunctions and $\delta(x-x^{\prime})$. He showed that such objects are not legitimate functions within Hilbert space and argued instead for the use of spectral measures—the idea he later developed into the spectral theorem in his 1932 book Mathematische Grundlagen der Quantenmechanik [36, 37]. The novelty in 1927 was not the final spectral theorem but the recognition that the continuous spectrum can be handled rigorously without $\delta$-functions.
He rephrased both matrix and differential eigenvalue equations as the single operator equation $Hf=af$, where $f$ belongs to a Hilbert space $\Phi$, consisting of sequences or square-integrable functions satisfying (123) [36]. This abstraction marks the birth of the spectral viewpoint: all observables are linear operators on a Hilbert space.
Physicists such as Born, Jordan, and Dirac had the correct physical ideas—probabilities, amplitudes, unitary equivalence—but their mathematics was loose. Von Neumann’s work provided the necessary rigor: proving convergence, defining operator domains, and formalizing the scalar products and norms that underlie the theory.
In his 1932 book, von Neumann gave the first fully rigorous unification of wave and matrix mechanics. He began from the Riesz–Fischer isomorphism between the sequence space of matrix mechanics [37]:
$F_{Z}=\left\{\,x=(x_{\nu})\;\big|\;\sum_{\nu=1}^{\infty}|x_{\nu}|^{2}<\infty\,\right\},$ (125)
@@
$F_{\Omega}=L^{2}(\Omega)=\left\{\,\phi:\Omega\to\mathbb{C}\;\big|\;\int_{\Omega}|\phi(q)|^{2}\,dq<\infty\,\right\}.$ (126)
which are linearly isometric:
@@
Under this correspondence, the basic observables match. In the wave picture, position and momentum act by:
$q_{j},\qquad\frac{\hbar}{i}\,\frac{\partial}{\partial q_{j}},$ (128)
while in the sequence space, they correspond to matrices $Q_{j},P_{j}$. Both satisfy the canonical commutation relations [37]:
@@
and any Hamiltonian $H$ formed algebraically from them yields identical predictions in either representation.
Von Neumann then abstracts to define Hilbert space as a complete, separable complex inner-product space. He analyzes orthonormal sets, expansions:
@@
and completeness. He systematizes operators, their domains, adjoints $A^{\ast}$, and the distinction between bounded and unbounded operators. The stationary Schrödinger equation becomes the abstract eigenvalue problem $H\phi=\lambda\phi$ for a self-adjoint operator $H$.
He formulated probabilities as quadratic forms [37]:
@@
where $E_{A}(\Delta)$ is the spectral projection of $A$. For the position operator $Q$ on $L^{2}(\mathbb{R}^{3})$, this reduces to Born’s rule:
@@
Thus the abstract formula $(\psi,E_{A}(\Delta)\psi)$ gives a representation-independent foundation for Born’s probability law.
@@
In this way, von Neumann completed the transition from quantum mechanics as a calculational scheme to an axiomatized physical theory. States become Hilbert-space vectors (or density operators); observables become self-adjoint operators; probabilities arise from spectral measures. Matrix mechanics and wave mechanics appear as two unitarily equivalent representations of a single Hilbert-space theory.